Fermi National Laboratory
 

Accelerator Update for August 2, 2006 - August 4, 2006

 

Wednesday August 2

The day shift began with Operations monitoring store 4868 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) off due to problems with A:IB, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with NuMI and MiniBooNE taking beam.

At 8:57 AM, NuMI tripped off on a “Target and LCW flow” indication. Operations contacted an expert and with his permission, reset the trip.

Operation began preparing Recycler at 11:32 AM to stash antiprotons. The stash ended.

Commonwealth Edison asked Fermilab to reduce its power usage. The Run Coordinator lowered the MI duty cycle to 5 seconds at 11:55 AM.

Operations reported at 3:27 PM that a MI expert had been conducting slip stacking studies for most of the afternoon.

The Recycler began preparing to stash at 6:04 PM. The stash ended and Pbar resumed stacking at 6:45 PM.

At 6:49 PM, Operations reported a problem with a MI RF cavity (MIRF3). Its gap envelope was not playing out properly. At 8:03 PM, an expert had operators make adjustments to the filament. The expert said that the power amplifier (PA) was old and getting weak.

The Recycler began preparing to stash at 9:11 PM.

At 10:08 PM, a lightning strike tripped off the MI power supply loop and the beam permit. The Recycler quadrupole ramp corrector loop (R:QCLP) also tripped off. Half of the stash was lost. Everything reset.

At 10:29 PM, the TeV quenched probably due to the multiple lightning strikes around the lab. Operators sequenced the TeV to 150 GeV to protect against further strikes.

At 10:30 PM, four Booster RF stations (BRF2, 11, 13, and 18) tripped off for various reasons.

At 10:35 PM, a Booster quadrupole power supply (B:Q6) tripped off. Operators could not turn it on. They called in a power supply expert.

The stash ended at 10:46 PM.

Thursday August 3

The midnight shift began with Operations monitoring the stash, and with Pbar, NuMI, and MiniBooNE off due to work on the Booster.

The power supply expert replaced the supply for B:Q6 at 12:14 AM. Pbar resumed stacking. NuMI and MiniBooNE resumed taking beam.

Operations put the TeV through a dry squeeze at 12:53 AM.

TeV experts began a study at 1:23 AM.

The Recycler began preparing to stash at around 5 AM. The stash ended.

At 5:09 AM, a Booster vacuum pump tripped off and took down BRF7 & 8. Everything reset.

At 5:21 AM, a Booster beam monitor tripped off, probably due to the electrical storm. Operators prepared the Booster for an access and by 6:10 AM had the monitor replaced.

Stacking resumed at 6:42 AM.

Operations put the TeV into shot setup at 6:59 AM.

At 7:15 AM, the Recycler (R:CE521) bulk supplies tripped off on an LCW flow. An expert reported at 7:49 AM that the problem was with the flow meter and not the flow; they bypassed the impeller. At 7:53 AM, they turned the supply back on and the stash was lost. This ended the shot setup.

Operations reported at 8:27 AM that Pbar had resumed stacking.

At 8:49 AM, a Pbar expert reported that the target had stopped rotating. He requested that the intensity be reduced while he tried to make the target spin.

A power supply expert reported at 9:19 AM that he had finished repairs on the MI-52 Recycler bulk power supply and had turned it on.

Operations prepared CDF and D0 for an access at 9:42 AM.

At 10:29 AM, Operations began turning off for accesses into the TeV, Booster, MI and Recycler, Pbar, the MI-8 line, and the 400 MeV area.

Operations turned off NuMI and MiniBooNE at 10:47 AM to conserve energy.

All accelerator accesses were over by 3 PM.

Operations established beam to MiniBooNE at 4:22 PM.

Operations reported at 8:01 PM that they had beam in MI and NuMI, and Pbar resumed stacking.

Recycler experts began aperture scans at 8:16 PM

The D0 and CDF access ended. Operations put the TeV into a dry squeeze at 9:05 PM.

The Recycler began preparing to stash at 9:16 PM. The stash ended and Pbar resumed stacking at 9:51 PM.

TeV experts began a study period at 10:30 PM.

Friday August 4

The midnight shift began with Operations monitoring the stash, with the TeV in a dry squeeze, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:32 AM, the MI south anode power supply tripped on an interlock fault. Operations called experts and had the supply on ten minutes later.

The Recycler began preparing to stash at 4:19 AM.

Operations put the TeV into shot setup at 6:24 AM.

At 7:56 AM, the TeV beam aborted due to a sector B1 low energy lead flow failure.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

Machine Reports

Linac

Current LINAC Status

  • Running on H- Source
    • Utilizing turbo pump only
    • The 750 keV transfer line pressure is elevated resulting in an additional 6% drop in Linac transfer efficiency. However, Linac losses appear very low and Booster is ok with beam current.
  • Source is on but in standby
  • Beam Current ~ 33 mA
  • Loss Sum Average ~ 15

Downtime Statistics

  • 92.9% Availability
  • 11.8 Hours w/ 19 Occurrences

Major Downtime ( >30 min. )

  • 5.27 Hours – numerous trips of QPS105 (Bad card found)
  • 3.58 Hours – Klystron Debuncher (bad driver module found)
  • Bad master Timing Pulse Repeater module LRF4

Booster

Last Week

  • Stacking: 10-12 Turns ~ 3.8 to 4.4E12 per batch
  • MiniBooNE: 10-11 Turns
  • NuMI: 10-12 Turns ~ 3.8 to 4.4E12 per batch

Studies

  • Efficiency problems due to high temperatures & the H- Source
  • Transition timing, bunch length, and head-tail instability measurement

Downtime: 6 hours (due to glitch) & 6 hours due to an access

  • RF miscellaneous
  • Beam positions
  • 400 MeV area quadrupole power supply
  • GMPS trips
  • Access: repaired 4 RF stations and the GMPS thermocouples

Next Week

  • Continue to work on orbits – correctors running too hard
  • Transition work will continue
  • Work to bring efficiency back up

Main Injector

MI operation

  • Shot to the TeV.
    • Pbar coalesce eff. 82 - 88%
  • Shot to the RR
    • Energy mismatch from Acc. To MI
  • Stacking
    • Running with mixed mode $23 cycle.
    • Intensity @injection ~ 7.5E12
    • Intensity @ target ~7.1E12 (Acc. eff. ~ 94%)

BPM upgrade

  • MI10 & 50 electronics installation -complete
  • All BPM installation has been completed
  • Repaired 2 BPM combiner in the tunnel

BLM upgrade

  • Still investigating the noise issues
  • 12 BLMs were disconnected in the tunnel because of a bad HV cable

MI studies

  • Multi-batch slip stacking study – next week

Pbar

Downtime / Failures this week

  • Pulsed Magnet (PMAG) failure.
    • On 7/29 PMAG ground faulted. When the replacement was done, measurements indicated the magnet was open, but it is not clear yet where. Closer inspection will be done when the magnet cools down. Debris was found inside, but was not seriously occluding the beam aperture.
  • D:IB (Debuncher dipole supply) ground faulted
    • Resistors in the filter inside the supply failed. Lucky for the tech, he was already here working on the PMAG supply
  • A:IB (Accumulator dipole supply) ground faulted
    • A ground-faulted phase between the transformer and supply was the problem. A spare phase in another conduit was used for now, but the correct fix is to pull 3 new phases through the same piece of conduit
  • AP0 target rotation motor replaced
    • Was done under MI access, so no lost stacking time

Machine work this week

  • P1-P2-AP1 tuning program (‘Overthruster’) continues to have problems with AP1 BPMs. We also have similar problems with AP3 BPMs when doing reverse proton tune-ups before transfers
  • Two adjustments were done to correct the energy match between Accumulator & MI on Recycler transfers. These energy changes are due to the hot weather (LCW temperature)
  • We’ve made more effort this past week to test and certify software for LINUX migration
  • Transfer efficiency problems from the Accumulator to the Recycler have been looked at. Some of the trouble is in how we save / restore beamline settings and is a sequencer problem we’re working on
  • Stack rate / Production efficiency problems continue to plague us. The hot weather is not helping as energy changes due to temperature are faster than the empirical way the Accumulator stacktail system is typically tuned

TeV

Store Summary

Store

Avg Init. Luminosity (E30)

Avg Deliv’d

Luminosity

(nb -1)

Time (hr)

How did it end?

Comments

4880

     

Quench

Coupling changes on ramp, D0 vertical position/angle changes.

4868

120.2

4392

25.3

Quench

Quench at E11U due to lightning strike.

4862

158.7

8337

53.8

Intentional

Lost 10% p-bars on ramp, tune changes on ramp.

4859

171.21

8445

45.8

Intentional

1 st average delivered luminosity, 2 nd average initial luminosity , D48H separator would not switch polarity.

Machine Maintenance and Studies

  • Measured coupling on the ramp
  • Measured the injection helix using page P163
  • Measured chromaticity w/O2 octupoles off @ 150GeV
  • Changed vertical position and angle at D0
  • Preliminary test of the E48 bumper system as an abort gap cleaner
  • Corrections to coupling on the ramp
  • Took data with C100

Problems

  • D48 separator would not switch polarity
  • F17K2 liquid level was low causing trip
  • E1 QPM crate replaced because of bad fan
  • F2 dump, 1900mF cap bank replaced
  • E11U quench due to lightning

Study Requests

  • Check to see if antiproton injection bump is closed
  • Bumper system as an abort gap cleaner
  • Orbit stabilization on ramp
  • TEL2 orbit and timing/tuning

Recycler

Antiprotons to the Recycler

  • 18 Shot setup, 98 transfers in to the Recycler
    • 932E10 antiprotons transferred into the Recycler
  • Last 5 transfers average efficiencies

Historical

Last 5 transfers

ACC to MI

95%

85%

Mi to RR

95%

90%

ACC to RR

93%

80%

Problems

  • ACC to MI eff
    • Energy matching between Accumulator and MI
    • Miss-synchronization of sequencers for fudge factors
  • MI to RR eff
    • Due to above problem large injection oscillation lead to larger emittances

Antiprotons to the TeV

  • 2 shot setup 18 parcels : 519.1 E10 transfer
    • 28.8 E10 Average Intensity per Parcel
    • Emittances
      • Transverse 5 pi
      • Longitudinal 1.5 eV sec
  • Average efficiencies
 

Historical (last20)

Last 2 shots

RR to MI

97%

97%

Mi to TeV

86%

83%

Highlights and Issues

  • DCCT
    • Fixed during Thursday’s access into MI tunnel
  • Flying wires
    • Disk problem → New computer
    • Timing problem → Fixed without having to make an access
  • Routine transfers in and out of the Recycler
  • Did loose ~230e10 this past week
    • 15e10 - During studies
    • 60e10 - MI buss glitch due to storm
    • 155e10 (whole stash) - Correction element chassis @ RR52 building

Study Summary

  • Continuing to look for ways to reduce emittance growth when mining
    • During shots
      • Increase transverse emittance before mining
      • More tune adjustments
      • Turn electron beam off
  • Alignment of the electron beam in the CS
    • Still in progress
      • We want to refine alignment of the magnetic field in the first cooling solenoid
  • Diffusion rate measurements
    • Goal: Compare diffusion rate of stochastically cooled and electron cooled antiprotons
    • Mishap !! Got an instability and lost 15e10
    • Saw interesting features in the process
      • Flat longitudinal distribution
      • Strong transverse cooling

Plans and Upcoming Studies

  • Tune space exploration
  • Change horizontal chromaticity
  • Emittance dilution
    • At injection with high number of antiprotons
  • Drag rate/ SA sanity checks
  • Cooling studies ?
  • Installation of an additional AC unit at MI-31
    • Outside work in progress
    • Choose convenient time to do inside work (i.e. punch holes in the wall)
      • Estimated time 8-12 hours
        • Long shutdown
        • Possibly just after a store is in the early morning (then the electron beam is not needed for quite a long time)


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/09/2006    email Fermilab